# SemRel2024: A Collection of Semantic Textual Relatedness Datasets for 13 Languages

### v4

@@ 3.1.2 Data curation
Similarly, we selected Spanish sentence pairs from semantic similarity datasets Agirre et al. (2014, 2015); Cer et al. (2017a), entailment datasets such as SICK-es Huertas-Tato et al. (2021) and NLI-es Araujo et al. (2022). Additionally, we sampled contiguous sentences from the summarisation dataset XL-Sum Hasan et al. (2021), DiscoEval Spanish Araujo et al. (2022), and questions of different types from Spanish QC Á. García Cumbreras et al. (2006).
@@
For Amharic, we paired sentences present in news articles from different Ethiopian news outlets Yimam et al. (2021). Similarly, the Hausa and Kinyarwanda datasets include pairs of sentences from news articles collected by Abdulmumin and Galadanci (2019) and Niyongabo et al. (2020), respectively. Sentences shorter than five words and longer than 20 were excluded and pairs were created using lexical overlap. Additionally, for Amharic, we excluded sentences with mixed languages to avoid confusing the annotators.
@@
For Indonesian, we collected sentences from Wikipedia texts present in the ROOTS split Laurençon et al. (2022); Setya and Mahendra (2018) and the IndoSum Kurniawan and Louvan (2018) datasets. IndoSum is a human-written summarization dataset consisting of pairs of news articles with their abstractive summaries. We parsed both corpora at a sentence-level and only selected sentences that are composed of five to fifteen words.
@@ 3.2 Data annotation and challenges
We inspected annotators with large disagreements to ensure the annotation procedure was correctly followed (i.e., their annotations made sense for native speakers). Very strong disagreements would serve as a red flag of poor data quality resulting in a more thorough review of the annotator quality. Hence, as a sanity check, we looked to see that sentences with high relatedness scores were more semantically related than those with low relatedness scores. The specific procedure for ensuring data quality depends on the annotation procedure of the team (e.g., those using AMT vs those who did not). Note that disagreements were not deleted, as they can serve as a useful signal for BWS. That is, for a tuple $\langle p_{0},p_{1},p_{2},p_{3}\rangle$, when annotators disagree on what is most related (e.g., $p_{1}$ or $p_{3}$), then it is an indication that $p_{1}$ and $p_{3}$ may be semantically close to each other. As all tuple annotations (twice the number of instances) are used to determine the final scores of the sentence pairs, this disagreement would lead to the two pairs ($p_{1}$ and $p_{3}$) getting scores that are close to each other. On the other hand, if a sentence pair consistently occurs in 4-tuples that have very low annotator agreement, then it is likely that the sentence pair is the source of disagreement. This can be due to various reasons such as the language use, code-switching, or the annotator’s familiarity with the topics discussed.
@@ 3.3 Postprocessing and data quality control
Finally, a team of native speakers manually spotchecked the scores to make sure that the relatedness scores made sense and to supplement the quantitative evaluation based on SHR.
@@ Appendix A Annotation
esp MuchoChine (MC) (9%), Spanish QC (4%), PAWS-X Paraphrasing dataset (16%), NLI-es (11%),
SICK-es (16%), STS (5%), SP (19.5%), XL-Sum (19.5%).

### v5

@@ 3.1.2 Data curation
Similarly, we select pairs of Spanish sentences from semantic similarity datasets such as STS Agirre et al. (2014, 2015); Cer et al. (2017a), entailment datasets such as SICK-es Huertas-Tato et al. (2021) and NLI-es Araujo et al. (2022), and paraphrasing datasets such as PAWS-X Yang et al. (2019). We also sampled contiguous sentences from XL-Sum Hasan et al. (2021) and BSO DiscoEval Spanish Araujo et al. (2022), and we included questions of different types from Spanish QC Á. García Cumbreras et al. (2006).
@@
For Amharic, we paired sentences present in news articles from different Ethiopian news outlets Yimam et al. (2021). Similarly, the Hausa and Kinyarwanda datasets include pairs of sentences from news articles collected by Abdulmumin and Galadanci (2019) and Niyongabo et al. (2020), respectively. Sentences shorter than five words and longer than 20 were excluded, and pairs were created using lexical overlap. Additionally, for Amharic, we excluded sentences with mixed languages to avoid confusing the annotators.
@@
For Indonesian, we collected sentences from Wikipedia texts present in the ROOTS split Laurençon et al. (2022); Setya and Mahendra (2018) and the IndoSum Kurniawan and Louvan (2018) datasets. IndoSum is a human-written summarization dataset consisting of pairs of news articles with abstractive summaries. We parsed both corpora at a sentence level and only selected sentences that were composed of five to fifteen words.
@@ 3.2 Data annotation and challenges
We inspected annotators with large disagreements to ensure the annotation procedure was correctly followed (i.e., their annotations made sense for native speakers). Very strong disagreements would serve as a red flag of poor data quality resulting in a more thorough review of the annotation quality. Hence, as a sanity check, we examined whether sentences with high relatedness scores were more semantically related than those with low relatedness scores. The specific procedure for ensuring data quality depends on the annotation procedure of the team (e.g., those using AMT vs those who did not). Note that disagreements were not deleted, as they can serve as a useful signal for BWS. That is, for a tuple $\langle p_{0},p_{1},p_{2},p_{3}\rangle$, when annotators disagree on what is most related (e.g., $p_{1}$ or $p_{3}$), then it is an indication that $p_{1}$ and $p_{3}$ may be semantically close to each other. As all tuple annotations (twice the number of instances) are used to determine the final scores of the sentence pairs, this disagreement would lead to the two pairs ($p_{1}$ and $p_{3}$) getting scores that are close to each other. On the other hand, if a sentence pair consistently occurs in 4-tuples that have very low annotator agreement, then it is likely that the sentence pair is the source of disagreement. This can be due to various reasons such as the language use, code-switching, or the annotator’s familiarity with the topics discussed.
@@ 3.3 Postprocessing and data quality control
Finally, a team of native speakers manually spot-checked the scores to make sure that the relatedness scores made sense and to supplement the quantitative evaluation based on SHR.
@@ Appendix A Annotation
esp MuchoChine (MC) (7%), Spanish QC (4%), PAWS-X (19.5%), NLI-es (12%), SICK-es (17%),
STS (5%), BSO (17.5%), XL-Sum (18%).